The heat of fusion of water is 333J/g. Determine the kilojoules of energy required to heat 95.2 grams of ice at -10.0 *C to the melting point and melt all of the ice. The specific heat of ice is 2.03 J/g *C.
Answer #1

Ti = -10.0 oC
Tf = 0.0 oC
here
Cs = 2.03 J/g.oC

Heat required to convert solid from -10.0 oC to 0.0 oC
Q1 = m*Cs*(Tf-Ti)
= 95.2 g * 2.03 J/g.oC *(0--10) oC
= 1932.56 J

Hf = 333.0 J/g

Heat required to convert solid to liquid at 0.0 oC
Q2 = m*Lf
= 95.2g *333.0 J/g
= 31701.6 J

Total heat required = Q1 + Q2
= 1932.56 J + 31701.6 J
= 33634.16 J

= 33.6 KJ
Answer: 33.6 KJ